OpenStudy (anonymous):

You invest $2,000 in an account that has an annual interest rate of 3.2%, compounded anually. How much money will be in the account after 12 years?

OpenStudy (anonymous):

Would be answer be $2,919? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

A = P(1+r/n)^(nt) A = 2000(1+0.032/1)^(1*12) A = 2,918.6792 A = 2,919 So you got it right again, nice job.

OpenStudy (anonymous):

Awesome! =)

jimthompson5910 (jim_thompson5910):

yes it's definitely awesome that this is starting to click for you